OpenStudy (anonymous):

write an equation of the line containing the given point and paralle to the given line. Express your answer in the form y=mx+b (8,9) x+4y=5

OpenStudy (anonymous):

\[y=-\frac{1}{4}x+\frac{5}{4}\]\[y-9=-\frac{1}{4}(x-8)\]\[y=-\frac{1}{4}x+11\]

OpenStudy (anonymous):

parallel have the same gradient x+4y=5 4y=-x+5 y=-x/4+5/4 slope =-1/4 (8,9) (x,y)=-1/4 y-9 / x-8 =-1/4 4(y-9)=-1(x-8) 4y-36=-x+8 4y=-x+44 y=-x/4+11